OpenStudy (anonymous):

if the legnth of the hypotenuse of a right triangle is 13. the length of the shorter leg is 7 less than the lenth of the smaller leg. How do i find the length of the shorter leg

OpenStudy (campbell_st):

well set up an equation let the smaller leg be x then the shorter leg is x - 7 so using pythaoras' theorem you have \[(x -7)^2 + x^2 = 13^2\] now you need to simplify then solve the equation.

OpenStudy (anonymous):

that would be x squared-49 plus x2 equals 13 squared right?

OpenStudy (campbell_st):

not quite you have a perfect square so its \[x^2 - 14x + 49 + x^2 = 169\]

OpenStudy (anonymous):

hmm i have 0 idea how to solve this

OpenStudy (campbell_st):

ok... so you get \[2x^2 - 14x - 120 = 0\] divide each term by 2 \[x^2 - 7x - 60 = 0\] do you know how to factor quadratics equations..?

OpenStudy (anonymous):

the smallest leg would be 5 correct?

OpenStudy (campbell_st):

correct.... its a pythagorean triad...

OpenStudy (anonymous):

could u also tell me this. it says the length of an equilateral triangle is 4. find he length of the altitude.

OpenStudy (campbell_st):

ok... so the altitude will be perpendicular to a side here is a reference site that explains altitude http://www.mathopenref.com/altitude.html the triangle looks like 

OpenStudy (anonymous):

what would the formula be?

OpenStudy (campbell_st):

its pythaoras' theorem again, you have a hypotenuse of 4 and short side of 2, find the other short side

OpenStudy (anonymous):

ok and another question is the side of one square is 3 inches longer than the side of a second square.if the sum of their areas is 130 inches squared, find the legnth of the side of each square.